# Galactic bars are already mature at Cosmic Noon: bar strength and flatness at $z\sim 1.5$

### v1

@@ Abstract
facilities: JWSTsoftware: astropy (astropy22), GALIGHT (ding22)
1 Introduction
@@
The galactic bar is a ubiquitous morphological feature, with the fraction of galaxies hosting a bar steadily increasing from $\sim 10-20\%$ at $z\sim 1$ to $\sim 70\%$ in the local Universe (e.g., sheth03; jogee04; elmegreen04; marinova07; sheth08; masters11; simmons14; rosas-guevara20; zhao20; leconte24; guo25; euclid25; geron25; salcedo25; pastras25). These elongated structures, consisting of stable stellar orbits spanning the inner regions of disk galaxies, arise from growing non-axisymmetric instabilities in dynamically cold disks, with a Toomre-$Q$ parameter $\approx 1.5-2.0$ (combes81; combes93; athanassoula13b). Once formed, these instabilities grow as stars lose angular momentum and join the elongated orbits making up the bar (semczuk24). Theoretical works suggest that dynamical friction with the dark matter halo (athanassoula02; athanassoula13) and high-angular-momentum gas accretion to the disk outskirts (bournaud05c) regulate bar strength. Tidal interactions have also been proposed to trigger bar formation and evolution (gerin90; noguchi96; miwa98; berentzen04; romanodiaz08; lokas14; martinezvalpuesta17; lokas18; peschken19; ansar25; zheng25). Attempts to verify this interaction-driven pathway observationally remain inconclusive in the local Universe (vandenbergh02; aguerri09; barway11; lee12). However, tidally induced bars may dominate at $z\gtrsim 2$ (bi22; tsukui24), although huang25 argue the opposite for a barred galaxy at $z\sim 2.5$.
@@
The growth of bars has also been characterized through changes in the surface brightness profiles along the major axis of the bar (gadotti11; anderson22; erwin23). The pioneering study of elmegreen85 divided bars into two profile types: “flat” and “exponential”. Parameterizing these profiles with Sérsic models, kim15 further subdivided flat bars into flat ($n<0.4$) and intermediate ($0.4<n<0.8$), using the Spitzer Survey of Stellar Structure in Galaxies (S4G; sheth10). Meanwhile, $n>0.8$ profiles are defined as exponential. Bars are expected to initially follow the exponential disk profile, and are therefore more common in low-mass, star-forming “late-type” systems (combes93; kruk18). In contrast, massive and evolved “early-type” systems host flat, mature bars (also called“peak” and “shoulder” profiles, with the shoulders resulting in the characteristic flatness; erwin23). $N$-body simulations show that as bars grow, particles are trapped in looped $x_{1}$ orbits, producing flat profiles with shoulders. These shoulders may first appear as temporary structures at $\sim 1-2\,\rm Gyr$ after bar formation, but become stable after $\gtrsim 2-3\,\rm Gyr$ (anderson22; beraldo23). However, noguchi96 suggested that shoulders, and hence flatness, can also result from tidal interactions. This aligns with studies indicating that tidally induced bars may prematurely evolve into mature states (miwa98; fragkoudi25; zheng25).
@@
The timing of local bar formation is linked to the epoch of “disk settling”, originally expected at $z\gtrsim 1$ (sheth08; kraljic12; kim15; kim21). However, the exact epoch remains unclear, with observations finding indications of bar-formation as early as $z\sim 3.5-4.0$ (smail23; constantin23; amvrosiadis25; geron25; guo25; leconte25). These may be lower limits as resolution effects would likely limit detection of early phases of bar formation. With bars now detected at such high redshifts, their nature becomes the key question: are they long-lived or transient structures? Do they exert a dynamical effect on the host disk comparable to that in the local Universe?
@@
This work explores an alternate way of answering these questions, by measuring the major-axis surface brightness profiles of bars at $z\sim 1.5$. We aim to determine if they are flat and “mature”, having persisted for $\gtrsim 1-2\,\rm Gyr$ (indicated by N-body simulations, without the inclusion of gas and star-formation; anderson22), and therefore since $z\sim 2.5-3.5$. Alternatively, if the bars exhibit exponential profiles, it would imply repeated destruction and reformation due to tidal disruptions and/or dissipation from high gas densities (e.g., elmegreen04; elmegreen05b; kraljic12). We also quantify their strengths and impact on their hosts by calculating the transverse force exerted by the bar potential, enabling a direct comparison of bar strengths at $z\sim 1.5$ with those in the local Universe (lee20).
2 Sample
@@
The goal of this study is the detection of bars, which are most consistently identified at $z>1$ in rest-frame near-IR wavelengths (guo23). For our sample, BK25 likewise found that stellar substructures such as spirals and bars are detected mainly in the F [^4] W image ($\sim 1.77\,\mu m$ rest-frame at $z\sim 1.5$). We therefore restrict our analysis to bars in the F [^4] W images, while a joint multi-wavelength analysis will be presented in future work.
3 Morphological Analysis
3.1 Ellipse fitting
@@
When observing the galactic disk face-on, stellar bars are non-axisymmetric structures with centers consistent with the host disk [^1] However, off-centered bars have also been observed (kruk17; deswardt15). Therefore, the higher ellipticity ($\equiv 1$-axis ratio) of the bar compared to the disk can be used to systematically identify the presence of a bar. The commonly used method involves fitting radially concentric ellipses to the flux isophotes of the image, with the expectation that the ellipses with the highest ellipticity correspond to the bar (e.g., wozniak95; jogee04; marinova07; yu22; salcedo25). This procedure has now also been implemented at $z>1$, although visual inspection is still necessary to ensure proper bar identification (guo23; guo25; leconte24; leconte25).
@@
We implement this method on the BK25 sample, with two examples shown in Fig. 1. We begin with the F [^4] W image. Using the photutils [^2] https://photutils.readthedocs.io/ (photutils) package, we fit progressively larger ellipses to fainter isophotes (Fig. 1, left in each panel). We start at the center determined from the bulge-disk F [^4] W image morphological modeling in BK25, and a semi-major axis (SMA) of 2 pixels ($\sim$ F [^4] W PSF FWHM in pixels). After the first ellipse is fit to the isophote at SMA == 2 pixels, the code incrementally increases the SMA by a factor of $0.1$ and fits the next isophote. We allow the progressive ellipse centers to shift freely to enable the detection of offset bars. The ellipticity (Fig. 1,top right, in each panel) and position angle (PA; Fig. 1, bottom right, in each panel) of the series of ellipses can thus be plotted as a function of their SMA. A highly elliptical structure like a bar will appear as a local maxima in the ellipticity vs. SMA plot, beyond which the larger ellipses fit isophotes representing the more circular disk structure.
@@
We then locate the peak in the ellipticity vs. SMA plot (local maxima with an increase >0.1>0.1 in ellipticity; condition defined in marinova07) within the disk $r_{\mathrm{90}}$ (radius of $90\%$ flux), also from the bulge-disk modeling in BK25. If the peak occurs at SMA values << PSF FWHM/2, we reject it, as liang24 found bar detectability decreases drastically below this threshold. Given our small sample of 32 objects, we use visual inspection to ensure that ellipses corresponding to the peak ellipticity represent bars and not other disk substructures, such as spiral arms and clumps, which can be clearly identified. We find these substructures produce secondary spikes in ellipticity (e.g., ID 1749 and 1861, both shown in Fig. 1), as has also been noted in other $z>1$ studies (e.g., guo23).
@@
A second criteria is that of the PA being constant ($10^{\circ}$; kim21) across the length of the bar, since it is a structure with a fixed rotational velocity. However, to do this, we have to determine the inner radius of the bar (or at least a range of SMA datapoints clearly tracing the bar) to define a region over which the PA would remain constant. The current resolution and depth prevents us from satisfactorily doing that for most of the sample. However, in cases of (visually) strong bars, the peak ellipticity appears as a plateau with a constant PA throughout. In these cases, we invoke the PA criteria and define the bar length as the SMA value within this plateau that also corresponds to a decrease in the phase angle $10^{\circ}$ following the drop in ellipticity within $\sim 0.03^{\prime\prime}$ (the pixel size; see ID 89 in Fig. 1).
@@
3.2 Masking the spiral arms
The primary goal of this work is to model the profiles of the bars (Sec. 3.3 and Sec. 3.4). We find that the presence of spiral arms interferes with such analyses, as indicated by secondary spikes in the ellipse-fitting method (Sec. 3.1). The presence of spiral arms in the sample galaxies have already been commented on in BK25 and a subset of these have been analysed in kalita25c. In the previous section we find that these structures interfere with the ellipse-fitting method by increasing the ellipticity of isophotes. Therefore, in this section, we create masks for the spiral arms in preparation of the image-based morphological analysis of the bars (Sec. 3.3 and Sec. 3.4). To achieve this, we create 2D masks of the logarithmic spirals in these galaxies using a polar-coordinate wavelet decomposition method developed in kalita25c. This method involves de-projecting the galaxy using the disk axis ratio (from the BK25 bulge-disk modeling), followed by a wavelet decomposition in polar ($r,\theta$) coordinates. Using the m=26m=2-6 modes (signals that repeat 262-6 times over $\theta=0-2\pi$), a “spiral-like” image is constructed that contains the spiral structures, if present in a galaxy.
In kalita25c, this step was followed by fitting a spiral galaxy model using GALFIT (peng02; peng10). However, this was only possible for six out of 32 galaxies, mainly due to additional disk substructures complicating the fit. We have now developed a simpler tool to “detect” logarithmic spirals in the spiral-like image by fitting simple spiral models. The parametric function in polar co-ordinates for each logarithmic spiral arm can be defined as:
@@ 3.2 Masking the spiral arms
Here the additional factor of $\Delta$ (in radians) is introduced, which controls the starting phase angle of each spiral and is helpful while fitting multiple arms. The parameters $x_{c},y_{c}$ are the galactic center. Finally the ‘$D$’ parameter which can be values either 1-1 or $1$ controls whether a spiral arm propagates clockwise or anti-clockwise. With each arm path defined using Eq. 1, and with a thickness of 5 pixels ($\sim$ PSF size), we can generate multi-arm spiral ‘skeletons’ over a parameter space defined by the number of arms, $\Delta$ and $\phi$. These skeletons are fit to the aforementioned spiral-like image and we optimise the parameters by maximising the contrast between the flux within the skeletons and the rest of the galaxy. Once the final spiral skeleton is identified, we create a mask by thickening each arm by 9 pixels ($\sim 2\,\rm kpc$ at $z=1.5$, the approximate width of the arms in the sample; kalita25c). We then undo the de-projection to obtain the mask for the original galaxy image.
@@
Finally, we remove any part of the spiral arm mask that falls within the ellipse corresponding to the bar (Sec. 3.1). This may happen since the bar will be included in the spiral-like image we use to generate the masks, since it is also a structure dominant in the $m=2$ mode. However, we modify this step for two of the galaxies (IDs: 89 and 1147) which feature clearly visible strong bars. As discussed in the previous section, the ellipticity vs. SMA plot of such galaxies show a plateau with an inner peak as well as an outer peak. Although we use the outer peak to fix the bar ellipse, we rather use the inner peak to determine the minimum radius where we mask the spiral arms. We take this conservative approach since the flagging occurs only on one side of the bar in both cases, and hence there is still sufficient flux for the proper surface brightness modelling.
3.3 1-D major axis profile
@@
The profile of a galactic bar can be assessed using the flux distribution across the galaxy along the bar SMA (bar-major-axis profiles, mainly applied for galaxies in the local Universe; e.g., erwin23; gusev03). This is commonly used to determine whether the bar follows the exponential radial light profile of the disk [^3] Disks are expected to be exponential even at $z>1$ (elmegreen05b) (exponential bars) or exhibits the characteristic shoulders leading to a flatter appearance (flat–intermediate bars). Once a bar is detected in a galaxy in our sample, we generate similar bar-major-axis profiles by using the position angle of the ellipse corresponding to each bar (Sec. 3.1) to define a long slit-like aperture of thickness =5=5 pixels ($\sim$ PSF FWHM) along the bar.
@@
Figure 2: The 1D surface brightness profiles along the bar major axis of the galaxies in our sample (continued in Fig. 3). Also shown are the major-axis profiles of the 2D best fit bar models555the 1D profiles of the 2D bar models are generated through the same procedure that was used to generate the 1D profiles from the galaxy images in Sec. 3.3 (Sec. 3.4) for comparison. The x-axis gives the radial distance in kpc from the galaxy center. The red dashed lines denote the $R_{\mathrm{bar}}$, which is considered the approximate extent of the bar. Hence, the region within these lines contain the majority of the bulge and bar flux. In each case, we also provide the bar $n$-value from the 2D morphological fitting (Sec. 3.4), as well as the resulting classification into flat, intermediate and exponential bar profiles. In the five galaxies with flat profiles, the characteristic shoulders beyond the central bulge can be seen, giving the appearance of a “flattening” of the profile within $R_{\mathrm{bar}}$. The gaps in the profile are regions which have been flagged as they likely are regions influenced by spiral arms (Sec. 3.2).
@@
Collapsing the slit aperture over its width yields the 1-D flux distribution, which contains contributions from the bulge, bar, and disk, with spiral arms already masked out. The resulting major-axis profiles, shown in Fig. 5, are then fit with a composite model consisting of three Sérsic (F [^4] W PSF-convolved) components: bulge ($n=4$), disk ($n=1$), and bar (n=0.11.5n=0.1-1.5). Details of the PSF generation is mentioned in the next section. We decide on the Sérsic index range for the bar based on the local Universe study by kim15. We find it to be the largest study that implements fitting of Sérsic profiles to the surface brightness of bars. Hence we implement the same range in this study, not just for the 1D profile fitting, but also for the 2D fitting discussed next. The fit is first optimized using the Python package SCIFIT, followed by an MCMC $\chi^{2}$-optimization procedure. Although this 1-D morphological fitting is less robust than a 2-D approach—since it cannot fully exploit structural dissimilarities between the components—it still serves as a method of characterizing the distinct bar major-axis profiles observed visually.
3.4 2-D morphological fitting
@@
The final step of our analysis aims to fully decompose the surface brightness distribution of our galaxies into bulge, bar, and disk contributions (Fig. 4, 5). As in Sec. 3.2, we use spiral arm masks to minimize contamination. Hence, these are the dominant features appearing in the residuals. Following the procedure in BK25, we first perform a bulge–disk dual Sérsic fit with the Python-based package GALIGHT [^6] https://github.com/dartoon/galight (ding22), which implements the forward-modeling tool LENSTRONOMY [^7] https://github.com/lenstronomy/lenstronomy (birrer18; birrer21). This approach provides access to the full posterior distribution of each fitted parameter, with optimization carried out using the Particle Swarm Optimizer (kennedy95, PSO). We use the F [^4] W PSF generated using the software PSFEx (bertin11) on the full COSMOS-Web mosaic from which the galaxy images have been extracted. Furthermore, the error images are taken from the public COSMOS-Web database. These include the contributions of the corresponding weight images and the count-based Poisson noise.
@@
  • Bar: Variable n=0.11.5n=0.1-1.5 Sérsic model, covering the full range of bar profiles observed in the S4G sample at $z\sim 0$ (kim15). We assume this range also applies at $z\sim 1.5$. The bar phase is initialized from the ellipse-fitting results, while the initial $r_{e}$ is set to $2\times$ the bulge $r_{e}$.
@@
We check for dependence of our results on our choice of the bulge $n$-value. Hence, we repeat our analysis with bulge $n$ fixed to $2,2.5,3,3.5$ since a high fraction of local bulges can have $n<4$ (gao20). Using $n=1$ results in degeneracy with the disk component and requires additional constraints which we have tried to avoid. We find that the bar profile (discussed throughout this paper) does not get influenced by our choice of bulge index within the range n=24n=2-4. The measured bar parameters remain within the corresponding uncertainties.

### v2

@@ Abstract
facilities: JWSTsoftware: astropy (Astropy Collaboration et al., 2022), GALIGHT (X. Ding et al., 2022)
I Introduction
@@
The galactic bar is a ubiquitous morphological feature, with the fraction of galaxies hosting a bar steadily increasing from $\sim 10-20\%$ at $z\sim 1$ to $\sim 70\%$ in the local Universe (e.g., K. Sheth et al., 2003; S. Jogee et al., 2004; B. G. Elmegreen et al., 2004; I. Marinova & S. Jogee, 2007; K. Sheth et al., 2008; K. L. Masters et al., 2011; B. D. Simmons et al., 2014; Y. Rosas-Guevara et al., 2020; D. Zhao et al., 2020; Z. A. Le Conte et al., 2024; Y. Guo et al., 2025; Euclid Collaboration et al., 2025; T. Géron et al., 2025; J. M. Espejo Salcedo et al., 2025; S. Pastras et al., 2025). These elongated structures, consisting of stable stellar orbits spanning the inner regions of disk galaxies, arise from growing non-axisymmetric instabilities in dynamically cold disks, with a Toomre-$Q$ parameter $\approx 1.5-2.0$ (F. Combes & R. H. Sanders, 1981; F. Combes & B. G. Elmegreen, 1993; E. Athanassoula, 2013). Once formed, these instabilities grow as stars lose angular momentum and join the elongated orbits making up the bar (M. Semczuk et al., 2024). Theoretical works suggest that dynamical friction with the dark matter halo (E. Athanassoula, 2002; E. Athanassoula et al., 2013) and high-angular-momentum gas accretion to the disk outskirts (F. Bournaud et al., 2005b) regulate bar strength. Tidal interactions have also been proposed to trigger bar formation and evolution (M. Gerin et al., 1990; M. Noguchi, 1996; T. Miwa & M. Noguchi, 1998; I. Berentzen et al., 2004; E. Romano-Díaz et al., 2008; E. L. Łokas et al., 2014; I. Martinez-Valpuesta et al., 2017; E. L. Łokas, 2018; N. Peschken & E. L. Łokas, 2019; S. Ansar et al., 2025; Y. Zheng & J. Shen, 2025). Attempts to verify this interaction-driven pathway observationally remain inconclusive in the local Universe (S. van den Bergh, 2002; J. A. L. Aguerri et al., 2009; S. Barway et al., 2011; G.-H. Lee et al., 2012). However, tidally induced bars may dominate at $z\gtrsim 2$ (D. Bi et al., 2022; T. Tsukui et al., 2024), although S. Huang et al. (2025) argue the opposite for a barred galaxy at $z\sim 2.5$.
@@
The growth of bars has also been characterized through changes in the surface brightness profiles along the major axis of the bar (D. A. Gadotti, 2011; S. R. Anderson et al., 2022; P. Erwin et al., 2023). The pioneering study of B. G. Elmegreen & D. M. Elmegreen (1985) divided bars into two profile types: “flat” and “exponential”. Parameterizing these profiles with Sérsic models, T. Kim et al. (2015) further subdivided flat bars into flat ($n<0.4$) and intermediate ($0.4<n<0.8$), using the Spitzer Survey of Stellar Structure in Galaxies (S4G; K. Sheth et al., 2010). Meanwhile, $n>0.8$ profiles are defined as exponential. Bars are expected to initially follow the exponential disk profile, and are therefore more common in low-mass, star-forming “late-type” systems (F. Combes & B. G. Elmegreen, 1993; S. J. Kruk et al., 2018). In contrast, massive and evolved “early-type” systems host flat, mature bars (also called“peak” and “shoulder” profiles, with the shoulders resulting in the characteristic flatness; P. Erwin et al., 2023). $N$-body simulations show that as bars grow, particles are trapped in looped $x_{1}$ orbits, producing flat profiles with shoulders. These shoulders may first appear as temporary structures at $\sim 1-2\,\rm Gyr$ after bar formation, but become stable after $\gtrsim 2-3\,\rm Gyr$ (S. R. Anderson et al., 2022; L. Beraldo e Silva et al., 2023). However, M. Noguchi (1996) suggested that shoulders, and hence flatness, can also result from tidal interactions. This aligns with studies indicating that tidally induced bars may prematurely evolve into mature states (T. Miwa & M. Noguchi, 1998; F. Fragkoudi et al., 2025; Y. Zheng & J. Shen, 2025).
@@
The timing of local bar formation is linked to the epoch of “disk settling”, originally expected at $z\gtrsim 1$ (K. Sheth et al., 2008; K. Kraljic et al., 2012; T. Kim et al., 2015, 2021). However, the exact epoch remains unclear, with observations finding indications of bar-formation as early as $z\sim 3.5-4.0$ (I. Smail et al., 2023; L. Costantin et al., 2023; A. Amvrosiadis et al., 2025; T. Géron et al., 2025; Y. Guo et al., 2025; Z. A. Le Conte et al., 2025). These may be lower limits as resolution effects would likely limit detection of early phases of bar formation. With bars now detected at such high redshifts, their nature becomes the key question: are they long-lived or transient structures? Do they exert a dynamical effect on the host disk comparable to that in the local Universe?
@@
This work explores an alternate way of answering these questions, by measuring the major-axis surface brightness profiles of bars at $z\sim 1.5$. We aim to determine if they are flat and “mature”, having persisted for $\gtrsim 1-2\,\rm Gyr$ (indicated by N-body simulations, without the inclusion of gas and star-formation; S. R. Anderson et al., 2022), and therefore since $z\sim 2.5-3.5$. Alternatively, if the bars exhibit exponential profiles, it would imply repeated destruction and reformation due to tidal disruptions and/or dissipation from high gas densities (e.g., B. G. Elmegreen et al., 2004, 2005; K. Kraljic et al., 2012). We also quantify their strengths and impact on their hosts by calculating the transverse force exerted by the bar potential, enabling a direct comparison of bar strengths at $z\sim 1.5$ with those in the local Universe (Y. H. Lee et al., 2020).
II Sample
@@
The goal of this study is the detection of bars, which are most consistently identified at $z>1$ in rest-frame near-IR wavelengths (Y. Guo et al., 2023). For our sample, BK25 likewise found that stellar substructures such as spirals and bars are detected mainly in the F [^4] W image ($\sim 1.77\,\mu m$ rest-frame at $z\sim 1.5$). We therefore restrict our analysis to bars in the F [^4] W images, while a joint multi-wavelength analysis will be presented in future work.
III Morphological Analysis
In the following subsections, we discuss the set of analyses we undertake on the F [^4] W images to derive the structural properties of the galactic bars in our sample.
III.1 Ellipse fitting
@@
When observing the galactic disk face-on, stellar bars are non-axisymmetric structures with centers consistent with the host disk [^1] However, off-centered bars have also been observed (S. J. Kruk et al., 2017; B. de Swardt et al., 2015). Therefore, the higher ellipticity ($\equiv 1$-axis ratio) of the bar compared to the disk can be used to systematically identify the presence of a bar. The commonly used method involves fitting radially concentric ellipses to the flux isophotes of the image, with the expectation that the ellipses with the highest ellipticity correspond to the bar (e.g., H. Wozniak et al., 1995; S. Jogee et al., 2004; I. Marinova & S. Jogee, 2007; S.-Y. Yu et al., 2022; J. M. Espejo Salcedo et al., 2025). This procedure has now also been implemented at $z>1$, although visual inspection is still necessary to ensure proper bar identification (Y. Guo et al., 2023, 2025; Z. A. Le Conte et al., 2024, 2025).
@@
We implement this method on the BK25 sample, with two examples shown in Fig. 1. We begin with the F [^4] W image. Using the photutils [^2] https://photutils.readthedocs.io/ (L. Bradley et al., 2022) package, we fit progressively larger ellipses to fainter isophotes (Fig. 1, left in each panel). We start at the center determined from the bulge-disk F [^4] W image morphological modeling in BK25, and a semi-major axis (SMA) of 2 pixels ($\sim$ F [^4] W PSF FWHM in pixels). After the first ellipse is fit to the isophote at SMA == 2 pixels, the code incrementally increases the SMA by a factor of $0.1$ and fits the next isophote. We allow the progressive ellipse centers to shift freely to enable the detection of offset bars. The ellipticity (Fig. 1,top right, in each panel) and position angle (PA; Fig. 1, bottom right, in each panel) of the series of ellipses can thus be plotted as a function of their SMA. A highly elliptical structure like a bar will appear as a local maxima in the ellipticity vs. SMA plot, beyond which the larger ellipses fit isophotes representing the more circular disk structure.
@@
We then locate the peak in the ellipticity vs. SMA plot (local maxima with an increase >0.1>0.1 in ellipticity; condition defined in I. Marinova & S. Jogee, 2007) within the disk $r_{\mathrm{90}}$ (radius of $90\%$ flux), also from the bulge-disk modeling in BK25. If the peak occurs at SMA values << PSF FWHM/2, we reject it, as X. Liang et al. (2024) found bar detectability decreases drastically below this threshold. Given our small sample of 32 objects, we use visual inspection to ensure that ellipses corresponding to the peak ellipticity represent bars and not other disk substructures, such as spiral arms and clumps, which can be clearly identified. We find these substructures produce secondary spikes in ellipticity (e.g., ID 1749 and 1861, both shown in Fig. 1), as has also been noted in other $z>1$ studies (e.g., Y. Guo et al., 2023).
@@
A second criteria is that of the PA being constant ($10^{\circ}$; T. Kim et al., 2021) across the length of the bar, since it is a structure with a fixed rotational velocity. However, to do this, we have to determine the inner radius of the bar (or at least a range of SMA datapoints clearly tracing the bar) to define a region over which the PA would remain constant. The current resolution and depth prevents us from satisfactorily doing that for most of the sample. However, in cases of (visually) strong bars, the peak ellipticity appears as a plateau with a constant PA throughout. In these cases, we invoke the PA criteria and define the bar length as the SMA value within this plateau that also corresponds to a decrease in the phase angle $10^{\circ}$ following the drop in ellipticity within $\sim 0.03^{\prime\prime}$ (the pixel size; see ID 89 in Fig. 1).
@@
III.2 Masking the spiral arms
The primary goal of this work is to model the profiles of the bars (Sec. III.3 and Sec. III.4). We find that the presence of spiral arms interferes with such analyses, as indicated by secondary spikes in the ellipse-fitting method (Sec. III.1). The presence of spiral arms in the sample galaxies have already been commented on in BK25 and a subset of these have been analysed in B. S. Kalita et al. (2025c). In the previous section we find that these structures interfere with the ellipse-fitting method by increasing the ellipticity of isophotes. Therefore, in this section, we create masks for the spiral arms in preparation of the image-based morphological analysis of the bars (Sec. III.3 and Sec. III.4). To achieve this, we create 2D masks of the logarithmic spirals in these galaxies using a polar-coordinate wavelet decomposition method developed in B. S. Kalita et al. (2025c). This method involves de-projecting the galaxy using the disk axis ratio (from the BK25 bulge-disk modeling), followed by a wavelet decomposition in polar ($r,\theta$) coordinates. Using the m=26m=2-6 modes (signals that repeat 262-6 times over $\theta=0-2\pi$), a “spiral-like” image is constructed that contains the spiral structures, if present in a galaxy.
In B. S. Kalita et al. (2025c), this step was followed by fitting a spiral galaxy model using GALFIT (C. Y. Peng et al., 2002, 2010). However, this was only possible for six out of 32 galaxies, mainly due to additional disk substructures complicating the fit. We have now developed a simpler tool to “detect” logarithmic spirals in the spiral-like image by fitting simple spiral models. The parametric function in polar co-ordinates for each logarithmic spiral arm can be defined as:
@@ III.2 Masking the spiral arms
Here the additional factor of $\Delta$ (in radians) is introduced, which controls the starting phase angle of each spiral and is helpful while fitting multiple arms. The parameters $x_{c},y_{c}$ are the galactic center. Finally the ‘$D$’ parameter which can be values either 1-1 or $1$ controls whether a spiral arm propagates clockwise or anti-clockwise. With each arm path defined using Eq. 1, and with a thickness of 5 pixels ($\sim$ PSF size), we can generate multi-arm spiral ‘skeletons’ over a parameter space defined by the number of arms, $\Delta$ and $\phi$. These skeletons are fit to the aforementioned spiral-like image and we optimise the parameters by maximising the contrast between the flux within the skeletons and the rest of the galaxy. Once the final spiral skeleton is identified, we create a mask by thickening each arm by 9 pixels ($\sim 2\,\rm kpc$ at $z=1.5$, the approximate width of the arms in the sample; B. S. Kalita et al., 2025c). We then undo the de-projection to obtain the mask for the original galaxy image.
@@
Finally, we remove any part of the spiral arm mask that falls within the ellipse corresponding to the bar (Sec. III.1). This may happen since the bar will be included in the spiral-like image we use to generate the masks, since it is also a structure dominant in the $m=2$ mode. However, we modify this step for two of the galaxies (IDs: 89 and 1147) which feature clearly visible strong bars. As discussed in the previous section, the ellipticity vs. SMA plot of such galaxies show a plateau with an inner peak as well as an outer peak. Although we use the outer peak to fix the bar ellipse, we rather use the inner peak to determine the minimum radius where we mask the spiral arms. We take this conservative approach since the flagging occurs only on one side of the bar in both cases, and hence there is still sufficient flux for the proper surface brightness modelling.
III.3 1-D major axis profile
@@
The profile of a galactic bar can be assessed using the flux distribution across the galaxy along the bar SMA (bar-major-axis profiles, mainly applied for galaxies in the local Universe; e.g., P. Erwin et al., 2023; A. S. Gusev & M. G. Park, 2003). This is commonly used to determine whether the bar follows the exponential radial light profile of the disk [^3] Disks are expected to be exponential even at $z>1$ (B. G. Elmegreen et al., 2005) (exponential bars) or exhibits the characteristic shoulders leading to a flatter appearance (flat–intermediate bars). Once a bar is detected in a galaxy in our sample, we generate similar bar-major-axis profiles by using the position angle of the ellipse corresponding to each bar (Sec. III.1) to define a long slit-like aperture of thickness =5=5 pixels ($\sim$ PSF FWHM) along the bar.
@@
Figure 2: The 1D surface brightness profiles along the bar major axis of the galaxies in our sample (continued in Fig. 3). Also shown are the major-axis profiles of the 2D best fit bar models555the 1D profiles of the 2D bar models are generated through the same procedure that was used to generate the 1D profiles from the galaxy images in Sec. III.3 (Sec. III.4) for comparison. The x-axis gives the radial distance in kpc from the galaxy center. The red dashed lines denote the $R_{\mathrm{bar}}$, which is considered the approximate extent of the bar. Hence, the region within these lines contain the majority of the bulge and bar flux. In each case, we also provide the bar $n$-value from the 2D morphological fitting (Sec. III.4), as well as the resulting classification into flat, intermediate and exponential bar profiles. In the five galaxies with flat profiles, the characteristic shoulders beyond the central bulge can be seen, giving the appearance of a “flattening” of the profile within $R_{\mathrm{bar}}$. The gaps in the profile are regions which have been flagged as they likely are regions influenced by spiral arms (Sec. III.2).
@@
Collapsing the slit aperture over its width yields the 1-D flux distribution, which contains contributions from the bulge, bar, and disk, with spiral arms already masked out. The resulting major-axis profiles, shown in Fig. 5, are then fit with a composite model consisting of three Sérsic (F [^4] W PSF-convolved) components: bulge ($n=4$), disk ($n=1$), and bar (n=0.11.5n=0.1-1.5). Details of the PSF generation is mentioned in the next section. We decide on the Sérsic index range for the bar based on the local Universe study by T. Kim et al. (2015). We find it to be the largest study that implements fitting of Sérsic profiles to the surface brightness of bars. Hence we implement the same range in this study, not just for the 1D profile fitting, but also for the 2D fitting discussed next. The fit is first optimized using the Python package SCIFIT, followed by an MCMC $\chi^{2}$-optimization procedure. Although this 1-D morphological fitting is less robust than a 2-D approach—since it cannot fully exploit structural dissimilarities between the components—it still serves as a method of characterizing the distinct bar major-axis profiles observed visually.
III.4 2-D morphological fitting
@@
The final step of our analysis aims to fully decompose the surface brightness distribution of our galaxies into bulge, bar, and disk contributions (Fig. 4, 5). As in Sec. III.2, we use spiral arm masks to minimize contamination. Hence, these are the dominant features appearing in the residuals. Following the procedure in BK25, we first perform a bulge–disk dual Sérsic fit with the Python-based package GALIGHT [^6] https://github.com/dartoon/galight (X. Ding et al., 2022), which implements the forward-modeling tool LENSTRONOMY [^7] https://github.com/lenstronomy/lenstronomy (S. Birrer & A. Amara, 2018; S. Birrer et al., 2021). This approach provides access to the full posterior distribution of each fitted parameter, with optimization carried out using the Particle Swarm Optimizer (J. Kennedy & R. Eberhart, 1995, PSO). We use the F [^4] W PSF generated using the software PSFEx (E. Bertin, 2011) on the full COSMOS-Web mosaic from which the galaxy images have been extracted. Furthermore, the error images are taken from the public COSMOS-Web database. These include the contributions of the corresponding weight images and the count-based Poisson noise.
@@
  • Bar: Variable n=0.11.5n=0.1-1.5 Sérsic model, covering the full range of bar profiles observed in the S4G sample at $z\sim 0$ (T. Kim et al., 2015). We assume this range also applies at $z\sim 1.5$. The bar phase is initialized from the ellipse-fitting results, while the initial $R_{e}$ is set to $2\times$ the bulge $R_{e}$.
@@
We check for dependence of our results on our choice of the bulge $n$-value. Hence, we repeat our analysis with bulge $n$ fixed to $2,2.5,3,3.5$ since a high fraction of local bulges can have $n<4$ (H. Gao et al., 2020). Using $n=1$ results in degeneracy with the disk component and requires additional constraints which we have tried to avoid. We find that the bar profile (discussed throughout this paper) does not get influenced by our choice of bulge index within the range n=24n=2-4. The measured bar parameters remain within the corresponding uncertainties.